# An Exposition of Pathfinding Strategies Within Lightning Network Clients

### v1

@@ VII Evaluation
While the success rate, or payment reliability, remains a crucial metric in pathfinding algorithm design, it is not the sole consideration. Factors such as the fee incurred, the path length, and the total timelock or latency along the path also play significant roles. Consequently, we conducted a comparative analysis of LN client variants based on these metrics. The distribution of fee across transaction amount bins is illustrated in Figure 3. Additionally, the distribution of path lengths observed across routing client variants is represented in Figure 4.
Table IV presents a summary of metrics, encompassing fee, path length, and timelock, observed across routing clients. To ensure the robustness of our comparative analysis regarding these metrics, we exclusively considered transactions where routing succeeded in four or more routing client variants. In our analysis of the fee metric, we calculated the ratio of the fee to the transaction amount and determined the median across all transactions under consideration. Similarly, we computed the average path length and timelock across all transactions considered.
@@
Regarding these metrics, as shown in Table IV, Eclair_case3 demonstrates effective performance in terms of fee along the path, providing the lowest fee ratio compared to other clients. LND2, in addition to achieving a high success rate, yields paths with lower fees but relatively longer path lengths. LND1, on the other hand, offers paths with significantly high fees but shorter path lengths and lower latency. Among the routing clients, LDK exhibits the highest fee ratio. However, for larger payment amounts, as can be seen in Figure 3d, LDK offers the lowest fee ratio compared to all other clients. CLN emerges as the top performer in terms of total timelock, offering paths with the lowest latency for routing payment amounts. Eclair_case1 and Eclair_case2 perform the best in terms of path length, while Eclair_case3 exhibits the highest path length and latency.
TABLE III: Success rates among Lightning Network clients across transaction amount ranges for uniform distribution graph
Amount Bins LND1 LND2 CLN LDK Eclair_case1 Eclair_case2 Eclair_case3
$10^{0}-10^{1}$ 98.414 98.414 98.414 98.414 98.414 98.414 98.414
$10^{1}-10^{2}$ 98.226 98.226 98.226 98.226 98.226 98.226 98.226
$10^{2}-10^{3}$ 96.157 95.917 96.317 96.397 96.317 96.317 96.237
$10^{3}-10^{4}$ 87.440 89.440 87.680 89.680 88.640 88.640 89.600
$10^{4}-10^{5}$ 64.480 72.720 66.000 69.920 69.920 69.920 70.960
$10^{5}-10^{6}$ 39.440 57.840 46.480 45.440 50.640 50.640 55.920
$10^{6}-10^{7}$ 23.040 40.400 30.400 16.560 34.640 34.640 37.440
$10^{7}-10^{8}$ 18.480 29.120 22.480 9.360 26.160 26.160 27.040
@@
TABLE IV: Comparison of fee ratio, path length, and timelock metrics across Lightning Network Clients for uniform distribution graph
Attribute LND1 LND2 CLN LDK Eclair_case1 Eclair_case2 Eclair_case3
Fee 0.051% 0.016% 0.030% 0.070% 0.053% 0.053% 0.012%
Path Length 4.431 6.277 4.614 4.541 4.421 4.421 6.555
Timelock 195.499 258.426 187.738 200.857 196.053 196.053 270.158
TABLE V: Success rates among Lightning Network clients across transaction amount ranges for bimodal distribution graph
Amount Bins LND1 LND2 CLN LDK Eclair_case1 Eclair_case2 Eclair_case3
$10^{0}-10^{1}$ 97.871 97.950 97.792 97.792 97.792 97.792 97.871
$10^{1}-10^{2}$ 96.429 96.104 96.347 96.429 96.266 96.266 96.023
$10^{2}-10^{3}$ 92.560 92.560 92.400 92.880 92.640 92.640 93.120
$10^{3}-10^{4}$ 71.840 73.040 72.560 73.600 72.640 72.640 72.800
$10^{4}-10^{5}$ 44.960 50.480 44.240 50.160 49.040 49.040 48.880
$10^{5}-10^{6}$ 24.080 33.280 24.160 25.360 29.680 29.680 29.680
$10^{6}-10^{7}$ 16.560 18.720 16.480 10.960 19.920 19.920 18.640
$10^{7}-10^{8}$ 13.920 16.880 14.400 5.680 16.400 16.400 15.440
@@
TABLE VI: Comparison of fee ratio, path length, and timelock metrics across Lightning Network Clients for bimodal distribution graph
Attribute LND1 LND2 CLN LDK Eclair_case1 Eclair_case2 Eclair_case3
Fee 0.069% 0.020% 0.055% 0.164% 0.100% 0.100% 0.020%
Path Length 4.463 6.291 4.577 4.489 4.412 4.412 6.580
Timelock 198.387 264.593 190.894 203.937 198.792 198.792 276.539
TABLE VII: Success rates for amounts in range 10-100 sats across different node category combinations for source and receiver
Source-Receiver LND1 LND2 CLN LDK Eclair_case1 Eclair_case2 Eclair_case3
Well-Well 100 100 100 100 100 100 100
Well-Fair 100 100 100 100 100 100 100
Well-Poor 100 100 100 100 100 100 100
Fair-Well 100 100 100 100 100 100 100
Fair-Fair 100 100 100 100 100 100 100
Fair-Poor 99.394 99.394 99.394 99.394 99.394 99.394 99.394
Poor-Well 98.387 98.387 98.387 98.387 98.387 98.387 98.387
Poor-Fair 98.204 98.204 98.204 98.204 98.204 98.204 98.204
Poor-Poor 96.774 96.774 96.774 96.774 96.774 96.774 96.774
TABLE VIII: Success rates for amounts in the range of $10^{3}$-$10^{4}$ sats across different node category combinations for source and receiver
Source-Receiver LND1 LND2 CLN LDK Eclair_case1 Eclair_case2 Eclair_case3
Well-Well 97.60 96.80 96 99.20 96.80 96.80 96.80
Well-Fair 88.623 89.222 88.024 91.617 91.617 91.617 90.419
Well-Poor 90.4 90.4 92 90.4 91.2 91.2 88.8
Fair-Well 83.832 86.228 84.431 86.228 83.832 83.832 85.030
Fair-Fair 73.653 85.629 76.647 84.431 80.240 80.240 85.629
Fair-Poor 74.251 83.234 77.844 81.437 76.647 76.647 82.036
Poor-Well 89.6 89.6 89.6 89.6 90.4 90.4 90.4
Poor-Fair 80.838 91.617 86.826 90.419 84.431 84.431 91.018
Poor-Poor 89.6 88 88.8 89.6 89.6 89.6 86.4
TABLE IX: Success rates for amounts in the range of $10^{5}$-$10^{6}$ sats across different node category combinations for source and receiver
Source-Receiver LND1 LND2 CLN LDK Eclair_case1 Eclair_case2 Eclair_case3
Well-Well 40.80 64.80 50.40 46.40 53.60 53.60 60.80
Well-Fair 27.108 39.157 31.325 30.723 36.145 36.145 39.157
Well-Poor 48 62.4 52.8 51.2 61.6 61.6 64
Fair-Well 22.289 37.952 28.916 33.735 32.530 32.530 34.940
Fair-Fair 12.048 16.867 12.651 13.253 15.663 15.663 15.663
Fair-Poor 18.675 28.313 22.892 21.687 27.711 27.711 25.904
Poor-Well 45.6 60 51.2 43.2 56.8 56.8 53.6
Poor-Fair 18.072 27.108 23.494 27.108 27.108 27.108 26.506
Poor-Poor 28.8 40.8 34.4 33.6 33.6 33.6 40
We conducted similar simulations on the bimodal distribution graph, to examine the performance of the LN clients when channel liquidity is predominantly concentrated on one side. As shown in Table V, LND2 yields the highest success rate across multiple amount bins, specifically for large payment amounts. While we used the default value ($3\cdot 10^{8}$) for the liquidity broadening scale $\mathrm{s}$ in bimodal probability estimation (16), we also tested alternative values, including $\frac{\mathrm{cap}_{e}}{10}$ — the same scale used during channel liquidity sampling. However, LND2 did not produce better results with these alternative values compared to the default. Similar to the uniform distribution case Table III, LDK performs better for low and lower mid-range amount bins compared to higher amount bins. However, Eclair_case1 and Eclair_case2 deliver competitive success rates and outperform Eclair_case3, unlike in the uniform distribution case. LND1 still demonstrates lower success rates, while CLN shows the lowest success rate across multiple bins.
@@
In comparison to the uniform distribution case results in Table III, success rates in the bimodal distribution graph are significantly lower. This decline in success rates is attributed to the bimodal distribution of channel liquidity. As the payment amount increases, it becomes challenging to identify reliable routes for payment delivery due to the scarcity of available liquidity. Moreover, Table VI reveals a substantial increase in the fee ratio across all LN clients under the bimodal distribution scenario. Despite the bimodal distribution graph, Eclair_case3 still achieves the lowest fee ratio. In addition, Eclair_case1 and Eclair_case2 display the shortest path lengths, while CLN maintains the lowest overall timelock.
@@
As discussed earlier in this section, evaluating routing across different node categories is important. Therefore, we simulated transactions on the uniform distribution graph, by selecting nodes from well, fairly, and poorly connected node categories, resulting in a total of 9 experiments, each consisting of 1,000 transactions. These experiments covered all possible combinations of node categories by selecting source and destination nodes from the three connectivity levels. Payment amounts for each transaction were determined using the previously mentioned strategy.
@@
In all possible node category combinations for selecting source and receiver nodes, routing clients achieve varying success rates in different bins, as illustrated in Table VII, Table VIII and Table IX.
@@
Since we classify nodes based on both the number of channels and their total capacity, fairly connected nodes, with capacities ranging between $10^{4}$ and $10^{6}$, can only accommodate transaction amounts less than $10^{6}$, resulting in bins ranging from $1$ to $10^{6}$. In contrast, well and poorly connected nodes may have capacities exceeding $10^{6}$, enabling transaction amounts from bins ranging from $1$ to $10^{8}$. Therefore, to ensure a meaningful comparison of results, we compared success rates within each bin separately. From Table VII, it can be observed that for small payment amounts (10-1000 sats), the success rates are high across all node category combinations. However, notably, for combinations involving poorly connected nodes, the success rates are relatively lower.
@@
Overall from Table VIII and Table IX, across all routing clients, we observed that well-connected nodes exhibit high success rates, along with low fees, low delay, and short path lengths. Fairly connected nodes, on the other hand, provide lower success rates, charge moderate fees, with moderate delay and path length. In case of poorly connected nodes, the success rate surpasses that of fairly connected nodes. This is attributed to poorly connected nodes having substantial capacity, enabling successful payment routing despite having fewer channels. However, it should be noted that poorly connected nodes typically impose high fees, experience high delay, and have longer path lengths which are undesirable for efficient routing.
From the results of these connectivity level routing evaluation experiments, we found that LND2 consistently delivers high success rates across all node categories, while LND1 emerges as a poor performer, especially across all node category combinations in Table IX, compared to all other clients. LDK demonstrates moderate reliability, across most of the node combinations. In situations where nodes are poorly connected, CLN emerges as a strong contender due to its competitive performance. Eclair_case3 exhibits higher success rates for well-connected nodes, while Eclair_case1 and Eclair_case2record competitive success rates across all node combinations. Notably, even within these experiments, Eclair_case3 consistently provides paths with the lowest fees, though with the highest path length and timelock. Conversely, LDK yields paths with the highest fees for lower payment amounts. Eclair_case1 and Eclair_case2 offers shorter path lengths, whereas CLN yields paths with the lowest delay.
Analyzing the results from all our experiments and interpreting them using the weight functions from Section VI, several insights can be inferred. The high success rates exhibited by LND2 across various experiments can potentially be attributed to its adept modeling of success probabilities using bimodal estimators. To further improve the performance of LND2, it will be important to fine-tune the liquidity broadening scale $\mathrm{s}$ in (16).

### v2

@@ VII-B1 LN Graph with Uniform Balance Distribution
While the success rate, or payment reliability, remains a crucial metric in pathfinding algorithm design, it is not the sole consideration. Factors such as the fee incurred, the path length, and the total timelock or latency along the path also play significant roles. Consequently, we provide also alternative performance metrics of LN client variants in Table IV, which includes a fee ratio value, average path length, and average timelock, for which information across all amount bins of Table III is aggregated. To ensure a fair comparison, we exclusively considered transactions for which routing succeeded in five or more routing client variants. The fee ratio of Table IV is calculated as the median of the fraction of incurred fees and the transaction amount, with the median taken across all transactions under consideration. To understand the distribution of fees across transaction amount bins, we also report the median fee ratio for each bin in Table V.
@@
Among the considered LN clients, Eclair3 demonstrates effective overall performance in terms of fee along the path, providing the lowest fee ratio compared to other clients. LND-un, in addition to achieving a high success rate, yields paths with lower fees but relatively longer path lengths. LND-bm provides paths with lower fees, moderate path lengths, and significantly higher latency. LND-ap, on the other hand, offers paths with significantly higher fees but shorter path lengths and lower latency. Among the routing clients, LDK-un and LDK-bm provide paths with high fees, with LDK-bm exhibiting the highest fee ratio. However, for larger payment amounts, as seen in Table V, LDK-un and LDK-bm offer the lowest fee ratios compared to all other clients. CLN emerges as the top performer in terms of total timelock, offering paths with the lowest latency for routing payment amounts. Eclair1 and Eclair2 perform the best in terms of path length, while Eclair3 exhibits the highest path length and latency.
[…]
Amount Bins LND-ap LND-bm LND-un CLN LDK-un LDK-bm Eclair1 Eclair2 Eclair3
$10^{0}-10^{1}$ 98.103 98.103 98.103 98.103 98.024 98.024 98.103 98.103 98.103
$10^{1}-10^{2}$ 98.139 98.139 98.139 98.139 98.139 98.139 98.139 98.139 98.139
$10^{2}-10^{3}$ 97.598 97.838 97.678 97.518 97.838 97.758 97.598 97.598 97.838
$10^{3}-10^{4}$ 88.480 90.080 90.320 89.360 89.040 89.200 89.280 89.280 90.400
$10^{4}-10^{5}$ 69.280 72.800 74.800 71.040 68.480 69.040 72.400 72.400 74.240
$10^{5}-10^{6}$ 50.160 59.040 66.720 54.480 46.560 45.440 61.360 61.360 65.280
$10^{6}-10^{7}$ 33.200 37.040 47.680 36.720 22.960 21.680 44.240 44.240 45.040
$10^{7}-10^{8}$ 27.840 27.840 36.560 30.400 16.960 16.240 34.800 34.800 35.200
@@
TABLE IV: Comparison of performance metrics across LN clients, uniform balance distribution
Attribute LND-ap LND-bm LND-un CLN LDK-un LDK-bm Eclair1 Eclair2 Eclair3
Fee Ratio (median) 0.0500% 0.0200% 0.0142% 0.0271% 0.0550% 0.0564% 0.0503% 0.0503% 0.0116%
Path Length (avg.) 4.4505 5.4956 6.1855 4.6342 4.5783 4.5703 4.4242 4.4242 6.4419
Timelock (avg.) 196.2460 232.0691 252.3624 186.7123 200.0157 199.7607 195.5912 195.5912 263.9637
[…]
Amount Bins LND-ap LND-bm LND-un CLN LDK-un LDK-bm Eclair1 Eclair2 Eclair3
$10^{0}-10^{1}$ 0.3148 0.3005 0.3005 0.3498 10.3963 10.3802 0.6015 0.6015 0.3007
$10^{1}-10^{2}$ 0.3050 0.3001 0.3001 0.3318 1.2588 1.2722 0.6011 0.6011 0.3001
$10^{2}-10^{3}$ 0.2080 0.1277 0.1277 0.1957 0.3017 0.3026 0.2628 0.2628 0.1277
$10^{3}-10^{4}$ 0.0504 0.0180 0.0176 0.0302 0.0531 0.0551 0.0422 0.0422 0.0176
$10^{4}-10^{5}$ 0.0200 0.0091 0.0069 0.0083 0.0100 0.0107 0.0185 0.0185 0.0052
$10^{5}-10^{6}$ 0.0103 0.0109 0.0064 0.0035 0.0016 0.0016 0.0105 0.0105 0.0045
$10^{6}-10^{7}$ 0.0063 0.0097 0.0059 0.0055 0.0008 0.0007 0.0082 0.0082 0.0057
$10^{7}-10^{8}$ 0.0055 0.0055 0.0056 0.0055 0.0041 0.0041 0.0056 0.0056 0.0055
[…]
Amount Bins LND-ap $\underset{(s=3e5)}{\text{LND-bm}}$ $\underset{(s=\frac{cap}{10})}{\text{LND-bm}}$ LND-un CLN LDK-un LDK-bm Eclair1 Eclair2 Eclair3
$10^{0}-10^{1}$ 97.630 97.472 97.551 97.472 97.630 97.551 97.551 97.551 97.551 97.472
$10^{1}-10^{2}$ 98.543 98.462 98.543 98.462 98.462 98.219 98.219 98.462 98.462 98.543
$10^{2}-10^{3}$ 93.435 93.755 94.476 94.155 93.595 92.714 92.874 93.755 93.755 94.396
$10^{3}-10^{4}$ 78.080 79.360 81.600 80.240 78.160 76.720 77.760 79.520 79.520 79.520
$10^{4}-10^{5}$ 51.120 55.280 61.600 58.400 51.520 47.040 49.360 54.640 54.640 57.280
$10^{5}-10^{6}$ 34.720 36.320 45.600 40.560 33.520 21.840 23.120 38.560 38.560 37.680
$10^{6}-10^{7}$ 26.960 28.160 35.920 33.120 26.080 13.360 13.760 30.480 30.480 29.760
$10^{7}-10^{8}$ 21.360 21.200 22.480 20.800 18.480 11.520 11.200 21.840 21.840 19.520
@@
TABLE VII: Comparison of performance metrics across LN clients, bimodal balance distribution
Attribute LND-ap $\underset{(s=3e5)}{\text{LND-bm}}$ $\underset{(s=\frac{cap}{10})}{\text{LND-bm}}$ LND-un CLN LDK-un LDK-bm Eclair1 Eclair2 Eclair3
Fee Ratio (median) 0.0509% 0.0201% 0.0202% 0.0175% 0.0355% 0.0817% 0.0867% 0.0571% 0.0571% 0.0148%
Path Length (avg.) 4.4528 5.6160 6.2359 6.2138 4.6012 4.5122 4.5026 4.3929 4.3929 6.4737
Timelock (avg.) 196.8925 238.4878 259.2788 257.4595 187.9205 199.8414 199.9199 196.5492 196.5492 268.5613
[…]
Amount Bins LND-ap $\underset{(s=3e5)}{\text{LND-bm}}$ $\underset{(s=\frac{cap}{10})}{\text{LND-bm}}$ LND-un CLN LDK-un LDK-bm Eclair1 Eclair2 Eclair3
$10^{0}-10^{1}$ 0.3068 0.3000 0.3000 0.3000 0.3211 10.3301 10.3301 0.6729 0.6729 0.3000
$10^{1}-10^{2}$ 0.3050 0.3001 0.3001 0.3001 0.3560 1.3145 1.3265 0.7098 0.7098 0.3001
$10^{2}-10^{3}$ 0.2283 0.1190 0.1190 0.1190 0.2283 0.3007 0.3005 0.2756 0.2756 0.1190
$10^{3}-10^{4}$ 0.0457 0.0188 0.0200 0.0184 0.0310 0.0486 0.0500 0.0405 0.0405 0.0182
$10^{4}-10^{5}$ 0.0200 0.0080 0.0101 0.0054 0.0079 0.0098 0.0100 0.0192 0.0192 0.0044
$10^{5}-10^{6}$ 0.0128 0.0124 0.0094 0.0062 0.0031 0.0013 0.0013 0.0112 0.0112 0.0033
$10^{6}-10^{7}$ 0.0078 0.0100 0.0106 0.0063 0.0039 0.0005 0.0004 0.0108 0.0108 0.0057
$10^{7}-10^{8}$ 0.0056 0.0056 0.0100 0.0100 0.0055 0.0025 0.0025 0.0056 0.0056 0.0056
[…]
Bins s 3e5 $\frac{cap}{2}$ $\frac{cap}{4}$ $\frac{cap}{6}$ $\frac{cap}{8}$ $\frac{cap}{10}$ $\frac{cap}{12}$ $\frac{cap}{14}$ $\frac{cap}{50}$ $\frac{cap}{100}$ $\frac{cap}{10^{3}}$ $\frac{cap}{10^{4}}$ $\frac{cap}{10^{6}}$ $\frac{cap}{10^{8}}$
$10^{4}-10^{5}$ 55.44 58.32 59.44 59.92 60.80 60.88 61.36 61.60 63.6 63.12 57.28 53.6 52.16 52.16
$10^{5}-10^{6}$ 38.88 46.80 47.20 47.68 48.56 50.08 50.40 50 47.52 44.40 37.20 36.16 36.16 36.16
$10^{6}-10^{7}$ 28.16 31.28 32.56 34 34.16 34.56 35.36 35.44 31.52 29.60 26.96 26.56 26.48 26.48
[…]
Source-Receiver LND-ap LND-bm LND-un CLN LDK-un LDK-bm Eclair1 Eclair2 Eclair3
Well-Well 100 100 100 100 100 100 100 100 100
Well-Fair 100 100 100 100 99.394 99.394 99.394 99.394 100
Well-Poor 100 100 100 100 100 100 100 100 100
Fair-Well 100 100 100 100 100 100 100 100 100
Fair-Fair 100 100 100 100 100 100 100 100 100
Fair-Poor 98.193 98.795 98.795 98.795 98.795 98.795 98.795 98.795 98.795
Poor-Well 100 100 100 100 100 100 100 100 100
Poor-Fair 95.783 96.988 96.988 96.386 96.988 96.988 96.988 96.988 97.590
Poor-Poor 99.187 99.187 99.187 99.187 99.187 99.187 99.187 99.187 99.187
[…]
Source-Receiver LND-ap LND-bm LND-un CLN LDK-un LDK-bm Eclair1 Eclair2 Eclair3
Well-Well 99.2 99.2 99.2 100 99.2 99.2 100 100 99.2
Well-Fair 89.820 94.012 94.611 92.216 91.018 90.419 93.413 93.413 94.012
Well-Poor 90.4 91.2 89.6 90.4 88.8 89.6 91.2 91.2 88.8
Fair-Well 99.401 97.605 95.210 97.605 89.820 91.617 99.401 99.401 95.210
Fair-Fair 85.629 92.814 94.012 87.425 86.826 88.623 86.228 86.228 92.814
Fair-Poor 88.623 92.814 93.413 89.820 83.234 83.832 88.024 88.024 92.814
Poor-Well 93.6 93.6 93.6 93.6 92.8 92 93.6 93.6 94.4
Poor-Fair 81.437 88.024 88.623 83.832 86.826 89.820 85.629 85.629 88.623
Poor-Poor 83.2 84.8 86.4 85.6 84 84 84.8 84.8 86.4
[…]
Source-Receiver LND-ap LND-bm LND-un CLN LDK-un LDK-bm Eclair1 Eclair2 Eclair3
Well-Well 64 73.6 77.6 67.2 52.8 50.4 74.4 74.4 75.2
Well-Fair 27.711 31.928 31.325 27.108 24.699 25.904 31.928 31.928 31.325
Well-Poor 56 57.6 63.2 56.8 51.2 48 59.2 59.2 60
Fair-Well 31.325 33.133 32.530 33.133 24.699 23.494 33.735 33.735 33.735
Fair-Fair 14.458 24.096 23.494 17.470 18.072 15.060 23.494 23.494 25.301
Fair-Poor 33.735 37.349 37.952 33.133 21.084 21.084 36.145 36.145 34.337
Poor-Well 46.4 54.4 59.2 52.8 36.8 36.8 57.6 57.6 60
Poor-Fair 30.120 34.337 37.349 27.108 24.096 25.301 34.337 34.337 31.325
Poor-Poor 36 43.2 46.4 40 30.4 31.2 47.2 47.2 46.4
VII-B2 LN Graph with Bimodal Balance Distribution
To obtain a more complete picture of the LN clients’ performance given channel balances that are possibly more realistic in the real LN graph, we report the results of a similar simulation on the bimodal distribution graph in Tables VI, VII and VIII, in which the channel liquidities are sampled from (42) with $s=\frac{\mathrm{cap}_{e}}{10}$, so that the liquidities are predominantly concentrated on either side of the channel.
In addition to LND-bm, which uses the default value of ${3\cdot 10^{5}\;\mathrm{sats}}$ for the liquidity broadening scale $\mathrm{s}$ in (19), we also include simulations for LND-bm with $\mathrm{s}$ set to $\frac{\mathrm{cap_{e}}}{10}$ in (19), the same scale we used for sampling channel balances in the bimodal distribution graph.
We observe in Table VI, that for an LN graph with bimodal balance distribution, LND-bm with broadening scale parameter $\mathrm{s}=\frac{\mathrm{cap_{e}}}{10}$ yields the highest success rates across all amount bins except for micropayments between $1$ and $10$ sats. In contrast, LND-bm with the default value of $\mathrm{s}$ shows comparatively lower success rates. Similar to the uniform distribution case, LND-ap and the default LND-bm exhibit similar performance levels. LND-un performs competitively, achieving the second-highest success rates among LN clients. CLN performs moderately, while LDK-un and LDK-bm deliver the lowest success rates across the amount bins. Eclair1 and Eclair2 demonstrate competitive success rates, slightly outperforming Eclair3, in contrast to their performance in the uniform distribution case.
@@ VII-B2 LN Graph with Bimodal Balance Distribution
In comparison to the uniform distribution case results of Table III, success rates in the bimodal distribution graph are significantly lower. We attribute this decline in success rates to the bimodal distribution of channel liquidity, which inherently limits the routing ability of a given channel in one of the two directions unless the payment amount is very small, unlike in the uniform distribution case. As the payment amount increases, it becomes challenging to identify reliable routes for payment delivery due to the scarcity of available liquidity.
However, Table VII shows only a slight increase in the fee ratio across all LN clients under the bimodal distribution scenario. Even in the presence of the bimodal distribution graph, Eclair3 continues to achieve the lowest fee ratio. In addition, Eclair1 and Eclair2 display the shortest path lengths, while CLN maintains the lowest overall timelock. Table VIII implies that LDK-un and LDK-bm, despite their high overall fee ratios, provide paths with the lowest fees for large payment amounts, similar to the uniform distribution case.
VII-B3 Choice of Broadening Scale $s$ in LND-bm
Since enhanced success rates are observed for the bimodal distribution graph when using $\frac{\mathrm{cap}_{e}}{10}$ for $\mathrm{s}$ in (19)—the same scale used during channel liquidity sampling—compared to the default value, we conduct an ablation study to further investigate this effect, which we report in Table IX. Specifically, 3750 transactions are simulated across bins ranging from $10^{4}$ to $10^{7}$, with $\mathrm{s}$ values varied as shown in Table IX. The simulation results reveal that the success rates improve when $\mathrm{s}$ aligns closely with the scale used for channel liquidity sampling. These findings suggest that tuning the liquidity broadening scale $\mathrm{s}$ in (19) based on historical channel liquidity data could significantly enhance payment reliability.
VII-B4 Effect of Node Connectivity Levels
@@
To study the effect of node connectivity levels in routing payments, we simulate transactions on the uniform distribution graph, by selecting nodes from well, fairly, and poorly connected node categories, resulting in a total of 9 experiments, each consisting of 1,000 transactions. These experiments cover all possible combinations of node categories by selecting source and destination nodes from the three connectivity levels.
@@
In all possible node category combinations for selecting source and receiver nodes, routing clients achieve varying success rates in different bins, as illustrated in Table X, Table XI and Table XII.
@@
Since we classify nodes based on both the number of channels and their total capacity, fairly connected nodes, with capacities ranging between $10^{4}$ and $10^{6}$, can only accommodate transaction amounts less than $10^{6}$, resulting in bins ranging from $1$ to $10^{6}$. In contrast, well and poorly connected nodes may have capacities exceeding $10^{6}$, enabling transaction amounts from bins ranging from $1$ to $10^{8}$. Therefore, to ensure a meaningful comparison of results, we compare success rates within each bin ranging from $1$ to $10^{6}$ separately. From Table X, we observe that for small payment amounts (10-100 sats), the success rates are high across all node category combinations. However, notably, for combinations involving poorly connected nodes, the success rates are relatively lower.
@@
Overall from Table XI and Table XII, across all routing clients, we observe that well-connected nodes exhibit high success rates, along with low fees, low delay, and short path lengths. Fairly connected nodes, on the other hand, provide lower success rates, charge moderate fees, with moderate delay and path length. In case of poorly connected nodes, the success rate surpasses that of fairly connected nodes. This is attributed to poorly connected nodes having substantial capacity, enabling successful payment routing despite having fewer channels. However, it should be noted that poorly connected nodes typically impose high fees, experience high delay, and have longer path lengths which are undesirable for efficient routing.
From the results of these connectivity level routing evaluation experiments, we find that LND-un delivers the highest success rates across multiple node categories, while LND-ap performs poorly when the receiving node is fairly connected. LDK-un and LDK-bm demonstrate low reliability across most node combinations. CLN exhibits moderate reliability; however, in situations where nodes are poorly connected, CLN performs competitively. Eclair1, Eclair2, and Eclair3 emerge as strong contenders due to their high reliability rates across various node categories.
VII-C Analysis of Results
Analyzing the results from all our experiments and interpreting them using the weight functions from Section VI, several insights can be inferred. LND-un, formulated through a modification of LND’s channel success probability estimate, results in payment paths with high success rates. While the incorporation of more complex models that also leverage historical or external information about channel balances is beyond the scope of our study, we think that this indicates that reliability can be further improved by an adept modeling of success probabilities in a LN client’s weight function.
Despite the relatively realistic modeling of success probabilities in LND-bm, we observed that LND-bm exhibits relatively moderate reliability with default constants in the weight function. Our ablation study of Table VIII suggests that success rates of LND-bm can potentially be improved by fine-tuning the liquidity broadening scale $\mathrm{s}$ in (19).